Find the periodic payment R required to accumulate a sum of S dollars over t years with interest earned at the rate of r%/year compounded m times a year. (Round your answer to the nearest cent.) S = 45,000, r = 6, t = 9, m = 2
Answer #1

Interest is compounded 2 times per year, which means it is half yearly compounding

So using PMT function in Excel we have

Periodic payment = =PMT(6%2,18,,45000) = $ 1,921.89 every half year

Here is the proof

Period for interest accumulation Amount FV Factor Future Value
(1 + 6%/2)^Period
17 $1,921.89                         1.65 $   3,176.59
16 $1,921.89                         1.60 $   3,084.07
15 $1,921.89                         1.56 $   2,994.24
14 $1,921.89                         1.51 $   2,907.03
13 $1,921.89                         1.47 $   2,822.36
12 $1,921.89                         1.43 $   2,740.16
11 $1,921.89                         1.38 $   2,660.35
10 $1,921.89                         1.34 $   2,582.86
9 $1,921.89                         1.30 $   2,507.63
8 $1,921.89                         1.27 $   2,434.59
7 $1,921.89                         1.23 $   2,363.68
6 $1,921.89                         1.19 $   2,294.84
5 $1,921.89                         1.16 $   2,228.00
4 $1,921.89                         1.13 $   2,163.11
3 $1,921.89                         1.09 $   2,100.10
2 $1,921.89                         1.06 $   2,038.93
1 $1,921.89                         1.03 $   1,979.55
0 $1,921.89                         1.00 $   1,921.89
$ 45,000.00
Answer #2

SOLUTION :


Following formula will be applicable for periodic payment $R to accumulate $S in t years at interest rate of r% per year compounded m times a year ;


S = R((1 + r / (100m))^(mt) - 1) / (r / (100m))   (ANSWER).


For S = $45000 ; r = 6% ; t = 9 years ; m = 2 , we have :


=> 45000 = R ( (1 + 6 / 200)^(18) - 1 ) / (6 / 200)

=> 45000 = R * 23.4144354

=> R = 45000/23.4144354

=> R = 1921.89 ($) (ANSWER).

answered by: Tulsiram Garg